OpenStudy (anonymous):

Make each number sentence true by inserting parentheses. 100 = 15+ 10 * 4 4 = 24 / 4 + 2

OpenStudy (apoorvk):

100 = ( 15+ 10) * 4 and 4 = 24 / (4 + 2)

OpenStudy (anonymous):

THANK YOU !!!!!

OpenStudy (anonymous):

Will you help me with the next few ones.

OpenStudy (anonymous):

8 = 24 / 4 = 2

OpenStudy (anonymous):

Hello

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

8 = 24 / 4 = 2

OpenStudy (anonymous):

Where do the parentheses go?

OpenStudy (apoorvk):

okay wait. you mean these are two equalities? and both to be solved separately?